OpenStudy (anonymous):

5x+7=22. I'm having a hard time. Can you give me the work and answer? Help me please. I'm really bad at math.

OpenStudy (anonymous):

subtract 7 divide by 5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

You should be able to do this on your own.

OpenStudy (anonymous):

Well I can't. I thought this was a website where you help people with math.

OpenStudy (skullpatrol):

I'll try and help you if you try and learn.

OpenStudy (anonymous):

I'm tryint=g and I still don't get it.

OpenStudy (anonymous):

Trying*

OpenStudy (skullpatrol):

Do you know what an inverse operation is?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

5x+7=22 -7 5x=15 x=3

OpenStudy (skullpatrol):

What grade are you in?

OpenStudy (anonymous):

8th.

OpenStudy (skullpatrol):

inverse operations: Operations that "undo" each other, for example multiplication and division.

OpenStudy (anonymous):

I literally don't get any of this.

OpenStudy (skullpatrol):

If the side containing the variable involves a certain order of operations apply the inverse operations in the opposite order.

OpenStudy (jiteshmeghwal9):

Subtract 7 both sides by transposition 5x+7-7=22-7 5x=15 Now divide 5 both sides by transposition 5x/5=15/5 what u gt as the value of x will be ur answer :)

OpenStudy (anonymous):

This is confusing. I'm not doing this. This website is worthless.

OpenStudy (skullpatrol):